OpenStudy (anonymous):

Could someone help me learn how to do this?? I will give a medal and fan to whoever help me learn the best! Solve the proportion 2x/5x + 6 = 2/x

OpenStudy (anonymous):

x = -1 and x = 6 x = -1 and x = -6 x = 1 and x = -5 x = 1 and x = 5

OpenStudy (anonymous):

@robz8

OpenStudy (anonymous):

so you're looking for \[\frac{ 2x }{ 5x+6 } = \frac{ 2 }{ x }\]?

hero (hero):

Okay, so it's \[\frac{2x}{5x+6} = \frac{2}{x}\]

OpenStudy (anonymous):

Okayy!

OpenStudy (anonymous):

cross multiply and solve for x

OpenStudy (anonymous):

Cross multiply?

OpenStudy (anonymous):

it's a way to get rid of the denominators by multiplying both sides of the equation by both of the denominators

OpenStudy (anonymous):

you end up with 2x^2 = 10x +12

OpenStudy (anonymous):

then move everything to one side 2x^2 - 10x - 12 divide everything by 2 x^2 - 5x - 6 factor (x-6)(x+1) which means x = 6 and -1 a faster way to answer that would just be to plug in the x values in the answers and see which pair worked :P

OpenStudy (anonymous):

Thank you as always!!